Situation: Lovely paid Php 600 for 10 kilos of rice and 3 kilos of fish.
a. What mathematical inequality represents the total amount Lovely paid? Define the variables used.
b. Suppose a kilo of rice costs Php 35. What could be the greatest cost of a kilo of fish to the nearest pesos?

Answer by CubeyThePenguin(3113)

a) r = price of 1 kilo of rice
f = price of 1 kilo of fish

10r + 3f <= 600

b) r = 35, so plug that in:

10(35) + 3f <= 600
350 + 3f <= 600
3f <= 250

f <= 250/3 or f <= 83 (rounded down to the nearest peso)